The diagonal of a square has the length of 12radical2 feet. Find the length of the sides of the square. Length of the side is ? feet....
Found 2 solutions by TimothyLamb, ABU1998:
Answer by TimothyLamb(4379)

diagonal for the general square:
h = sqrt( aa + aa )
h = sqrt( 2aa )
h = sqrt( 2 ) * sqrt( aa )
h = sqrt( 2 ) * a
---
this square:
h = sqrt( 2 ) * 12
---
equate h's:
sqrt( 2 ) * a = sqrt( 2 ) * 12
a = sqrt( 2 )/sqrt( 2 ) * 12
a = 12
---
answer:
side length = 12 feet
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by ABU1998(5)

the square has 4 equal sides,so we use the Pythagorean theorem:
hypotenuse=side square +side square so
h^2=s^2 +s^2=2s^2 ,substitute for the hypotenuse 12sqrt12
(12sqrt12)^2 = 2.s^2 ,this gives 144(12) = 2.s^2 ,divide by 2
144(6) = s^2 so take the square root of each side
12.sqrt(6) = s ,I hope it helps